OpenStudy (anonymous):

what is the sum of 3√5 and 2√20?

OpenStudy (zzr0ck3r):

\[hint:\space 2\sqrt{20}=2\sqrt{5*4}=2\sqrt{4}\sqrt{5}=2*2*\sqrt{5}=4\sqrt{5}\]

OpenStudy (zzr0ck3r):

so what is 3 of something + 4 of something?

OpenStudy (anonymous):

7√5,30√2,3√5+10√2,and 60 these are the choices

OpenStudy (anonymous):

3+sqrt(5), 4 sqrt(5)^^

OpenStudy (zzr0ck3r):

what is \[3\sqrt{5}+4\sqrt{5}?\]

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

*high five*

OpenStudy (anonymous):

still here?

OpenStudy (anonymous):

this is a summer school quiz so i need help with a couple more

OpenStudy (anonymous):

ok i guess ya gonna help..?

OpenStudy (anonymous):

.. lemme see one more

OpenStudy (anonymous):

how do you simplify √98x?

OpenStudy (anonymous):

and √x^16

OpenStudy (anonymous):

well for one the simplified version of sqrt of 98x would give you the results of \[7\sqrt{2} , \sqrt{x}\]